OpenStudy (mrnebula):

What is the length of the hypotenuse of a right triangle with legs of 9 cm and 12 cm.

OpenStudy (texaschic101):

Use the pythagorean theorem that states : a^2 + b^2 = c^2 where a and b are the sides and c is the hypotenuse. a^2 + b^2 = c^2 9^2 + 12^2 = c^2 81 + 144 = c^2 225 = c^2 square root of 225 = c 15 = c <-- length of hypotenuse is 15 cm.